OpenStudy (anonymous):

What are the roots of x in the equation -4 – 1`X`4` = x + 5?

OpenStudy (phi):

what does the ` mark mean ?

OpenStudy (anonymous):

\[What are the roots of x \in the equation -4 – \frac{ 14 }{ x } = x + 5?\]

OpenStudy (perl):

thats better :)

OpenStudy (phi):

to solve, first multiply the equation by x (that means both sides and *all terms*) can you do that ?

OpenStudy (mikurout):

the equation is x^2 + 9x + 14 = 0 sole it , then you will have your answers..